# SPECIAL RELATIVITY

### v9

@@ 15.6.3 On the Advanced ”Paradox” Related to the Coupling Fields and Particles
Now we consider another situation. Let us see what happens with a kicker, which is installed in the straight section upstream of the bending magnet and is characterized by a kick angle $\theta_{k}=\theta_{0}$. When the kick in the $y$ direction is introduced, there is a red shift of the critical wavelength which arise because, according to corrected addition velocities law, the electron velocity decreases from $v$ to $v-v\theta_{0}^{2}/2$ after the kick. The red shift of the critical frequency $\omega_{c}$ can be expressed by the formula $\Delta\omega_{c}/\omega_{c}=-(3/2)\gamma^{2}\theta_{0}^{2}$. We see a second order correction $\theta_{0}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$. The result of the covariant approach clearly depends on the absolute value of the kick angle $\theta_{0}$ and the radiation along the velocity direction has a red shift only when the kick angle has nonzero value. The implicit ”absolute” acceleration means acceleration relative to the fixed stars.
@@
The difference between these two situations, ending with a final uniform translation along the direction of the magnetic field is very interesting. It comes about as the result of the difference between two Lorentz coordinate systems in the lab frame. By trying to accelerate the electron upstream the bending magnet we have changed Lorentz coordinates for that particular source. We know that in order to keep a Lorentz coordinates system in the lab frame after the kick we need to perform a clock resynchronization. So we should expect the electron velocity to be changed. The difference between the two setups is understandable. When we do not perturb the electron motion (relative to the fixed stars) upstream of the bending magnet, no clock resynchronization takes place, while when we do perturb the motion, clock resynchronization is introduced.
@@
Figure 70: Two sets of initial conditions resulting in the same uniform motion along the magnetic field direction in the case of absolute time coordinatization in the lab frame. The magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups.
@@
We would now like to describe an apparent paradox. The argument that the difference between these two situations, ending with a final uniform translation along the magnetic fields direction, is paradoxical can be summarized in the following way: in the case of absolute time coordinatization in the lab frame, the initial conditions at the bending magnet entrance are apparently identical. In fact, the magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups. We must conclude that when we accelerate the electron in the lab frame upstream the bending magnet, the information about this acceleration is not included into the non-covariant trajectory. Where is the information about the electron acceleration recorded in the case of absolute time coordinatization? Since an electron is a structureless particle, the situation seems indeed paradoxical.
The above statement includes one delicate point. It is not true that an ultrarelativistic electron is a ”structureless particle”. Electrodynamics deals with observable quantities. Let us consider the measurement of the red shift in the bending magnet radiation from our kicked electron. We can measure the accurate value of the red shift using a spectrometer in the lab frame, and this leads to a description of the setup in the space-frequency domain.
@@
When the electron passes through a kicker, its fields are perturbed, and now include information about the acceleration. According to the old kinematics, the orientation of the virtual radiation phase front is unvaried. However, Maxwell’s equations do not remain invariant with respect to Galilean transformations and, as discuss throughout this book, the choice of the old kinematics implies using anisotropic field equations. As a result, the phase front remains plane but the direction of propagation is not perpendicular to the phase front. In other words, the radiation beam motion and the radiation phase front normal have different directions. Then, having chosen the absolute time synchronization, electrodynamics predicts that the virtual radiation beam propagates in the kicked direction with the phase front tilt $\theta_{k}$. This is the key to the ”paradox” discussed here. The information about the electron acceleration is recorded in the perturbation of the self-electromagnetic fields of the electron. Mathematically information is recorded in the phase front tilt of the virtual radiation beam.
15.7 Bibliography and Notes
1. It should be note that results of the beam splitting experiment at LCLS confirm our correction for spontaneous undulator emission [62]. It apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of the motion of the kicked beam. Therefore, coherent radiation from the undulator placed after the kicker is emitted along the kicked direction practically without suppression (see the Chapter 16 for more detail). In the framework of the conventional theory, there is a second outstanding puzzle concerning the beam splitting experiment at the LCLS. In accordance with conventional undulator radiation theory, if the modulated electron beam is at perfect resonance without kick, then after the kick the same modulated beam must be at perfect resonance in the velocity direction. However, experimental results clearly show that when the kick is introduced there is a red shift in the resonance wavelength. The maximum power of the coherent radiation is reached when the undulator is detuned to be resonant to the lower longitudinal velocity after the kick [62]. It should be remarked that any linear superposition of a given radiation field from single electrons conserves single-particle characteristics like parametric dependence on undulator parameters and polarization. Consider a modulated electron beam kicked by a weak dipole field before entering a downstream undulator. Radiation fields generated by this beam can be seen as a linear superposition of fields from individual electrons. Now experimental results clearly show that there is a red shift in the resonance wavelength for coherent undulator radiation when the kick is introduced. It follows that the undulator radiation from the single electron has red shift when the kick is introduced as well. This argument suggests that results of the beam splitting experiment in reference [62] confirm our correction for spontaneous undulator emission.
2. A widely accepted expression for the angular and spectral distributions of radiation from an ultra-relativistic electron on a helical orbit were calculated in [57, 58]. At present, relativistic cyclotron radiation results are textbook examples (see e.g. [59]) and do not require a detail description.
3. The covariant way of analyzing the radiation for helical motion was considered in [60]. It is generally believed that $\vec{x}(t)=\vec{x}(t)_{cov}$ and this is the reason why in the [60] there is no distinction between the two (non covariant and covariant) approaches to describe the electron motion on a helix downstream of the kicker setup.
@@ 16.1 Introductory Remarks
In the previous chapter we attempted to answer the question of why the error in radiation theory should have so long remained undetected. According to covariant approach, the various relativistic kinematics effects concerning to the synchrotron radiation setup, turn up in successive orders of approximation. Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$. In our previous discussion of bending magnet radiation, we learned that a motion of the single ultrarelativistic electron in a constant magnetic field, according to the theory of relativity, influences the kinematics terms of the second order $(v_{\perp}/c)^{2}$ only. It is demonstrated that due to a combination of the ultrarelativistic (i.e. paraxial) approximation and a very special symmetry of the conventional synchrotron radiation setup there is a cancellation of the second order relativistic kinematics effects except the non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion in the (bending) magnetic field direction. But synchrotron radiation from a bending magnets is emitted within a wide range of frequencies, and the output intencity is not sensitive on the red shift. That means that the spontaneous synchrotron radiation does not show sensitivity to the difference between covariant and non-covariant particle trajectories.
@@
But in the 21st century with the operation XFELs this situation changes. An XFEL is an example where the first order kinematics term $(v_{\perp}/c)$ plays an essential role in the description of the XFEL radiation and, in this case, covariant coupling of fields and particles predicts an effect in complete contrast to the conventional treatment. In this chapter we present a critical reexamination of existing XFEL theory. The main emphasize of this chapter is on coherent undulator radiation from the modulated electron beam. This chapter mainly addressed to readers with limiting knowledge of accelerator and XFEL physics.
@@
Figure 71: A well-known result of conventional (non-covariant) particle tracking. A micro-bunching electron beam passing through a weak dipole magnet (kicker) and undergoes a kick of an angle $\theta_{k}$. The propagation axis of the electron beam is deflected, while the wavefront orientation is preserved.
@@
The usual XFEL theory based on the use of old Newtonian kinematics for particle dynamics and the Einstein’s kinematics for the electrodynamics. In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities. For rectilinear motion of the modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking. However, one of the consequences of the relativity of simultaneity (i.e. mixture of positions and time) is a difference between covariant and non-covariant kinematics of a modulated electron beam in a given magnetic field. The theory of relativity shows that discussed above difference related with the acceleration along curved trajectories.
@@
There are several cases where the first order relativistic effect can occur in XFELs, mainly through the introduction of an trajectory kick (1). The most elementary of the effect that represents a crucial test of the correct coupling fields and particles is a problem involves the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
@@
It would be well to begin with bird’s view of some of the main results. Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion. As well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before (Fig. 71). In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e. in a wavefront tilt). In existing XFEL theory the wavefront tilt is considered as real. According to conventional treatment, a transverse kick does not change the orientation of a modulation wavefront, and hence suppresses the radiation emitted in the direction of the electron motion (2).
@@
Figure 72: A result of covariant particle tracking. In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e. the orientation of plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulation electron beam is treated using Lorentz coordinates. The theory of relativity dictates that a modulated electron beam in the ultrarelativistic limit has the same kinematics, in Lorentz coordinates, as laser beam. According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to the propagation direction.
@@
The covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment. Namely, in the ultrarelativistic limit, the wavefront of modulation, that is a plane of simultaneity, is always perpendicular to the electron beam velocity (Fig. 72). As a result, the Maxwell’s equations predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction. Experiments show that this prediction is, in fact, true (3). The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory (4).
@@
It is worth remarking that the absent of a dynamical explanation for the modulation wavefront readjusting in the Lorentz coordinatization has disturbed some XFEL experts. We only wish to emphasize that a good way to think of the modulation wavefront readjusting is to regard it as a result of transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
@@ 16.2 Modulation Wavefront Orientation
Let us suppose that a modulated electron beam moves along the $z$-axis of a Cartesian $(x,y,z)$ system in the lab frame. As an example, suppose that the modulation wavefront is perpendicular to the velocity $v_{z}$. How to measure this orientation? A moving electron bunch changes its position with time. The natural way to do this is to answer the question: when does each electron cross the $y$-axis of the reference system? If we have adopted a method for timing distant events (i.e. a synchronization convention), we can also specify a method for measuring the orientation of the modulation wavefront: if electrons located at the position with maximum density cross the $y$-axis simultaneously at certain position $z$, then the modulation wavefront is perpendicular to $z$-axis. In other words, the modulation wavefront is defined as a plane of simultaneous events (the events being the arrival of particles located at maximum density): in short, a ”plane of simultaneity”.
@@
Let us formulate the initial conditions in the lab frame in terms of orientation of the modulation wavefront and beam velocity. Suppose that $v_{z}$ is the velocity of the comoving frame $R(\tau)$ with respect to the lab frame $K(\tau)$ along the common $z$-axis in positive direction. In the lab frame we select a special type of coordinate system, a Lorentz coordinate system to be precise. Within a Lorentz frame (i.e. inertial frame with Lorentz coordinates), Einstein’s synchronization of distant clocks and Cartesian space coordinates $(x,y,z)$ are enforced. In order to have this, we impose that $R$ is connected to $K$ by the Lorentz boost $L(\vec{v}_{z})$, with $\vec{v}_{z}$, which transforms a given four vector event $X$ in space-time into $X_{R}=L(\vec{v_{z}})X$.
@@
We now consider the acceleration of the beam in the lab frame up to velocity $v_{y}$ along the $y$-axis. The question arises how to assign synchronization in the lab frame after the beam acceleration. Before acceleration we picked a Lorentz coordinate system. Then, after the acceleration, the beam velocity changes of an small value $v_{y}$ along the $y$-axis. Without changing synchronization in the lab frame after the particle acceleration we have a complicated situation as concerns electrodynamics of moving charges. As a result of such boost, the transformation of time and spatial coordinates has the form of a Galilean transformation. In order to keep a Lorentz coordinate system in the lab frame after acceleration, one needs to perform a clock resynchronization by introducing the time shift $t\to t+yv_{y}/c^{2}$. This form of time transformation is justified by the fact that we are dealing with a first order approximation. Therefore, $v_{y}/c$ is so small that $v_{y}^{2}/c^{2}$ can be neglected and one arrives at the coordinate transformation $y\to y+v_{y}t$, $t\to t+yv_{y}/c^{2}$. The Lorentz transformation just described differs from a Galilean transformation just by the inclusion of the relativity of simultaneity, which is only relativistic effect that appearing in the first order in $v_{y}/c$. The relation $X_{R}=L(\vec{v_{z}})L(\vec{v_{y}})X$ presents a step-by-step change from the lab reference frame $K(\tau+d\tau)$ to $K(\tau)$ and then to the proper reference frame $R$. The shift in the time when electrons located at the position with maximum density cross the $y$-axis of the lab frame $\Delta t=yv_{y}/c^{2}$ has the effect of a rotation the modulation wavefront on the angle $v_{z}\Delta t/y=v_{z}v_{y}/c^{2}$ in the first order approximation. In ultrarelativistic limits, $v_{z}\simeq c$, and the modulation wavefront rotates exactly as the velocity vector $\vec{v}$.
@@
What does this wavefront readjustment mean in terms of measurements? In the absolute time coordinatization the simultaneity of a pair of events has absolute character. The absolute character of the temporal coincidence of two events is a consequence of the absolute time synchronization convention. According to this old kinematics, the modulation wavefront remains unvaried. However, according to the covariant approach we establish a criterion for the simultaneity of events, which is based on the invariance of the speed of light. It is immediately understood that, as a result of the motion of electrons along the $y$ axis (i.e. along the plane of simultaneity before the boost) with the velocity $v_{y}$, the simultaneity of different events is no longer absolute, i.e. independent of the kick angle $\theta=v_{y}/c$. This reasoning is in analogy with Einstein’s train-embankment thought experiment.
@@
The wavefront orientation has no exact objective meaning, because the relativity of simultaneity takes place. The statement that the wavefront orientation has objective meaning to within a certain accuracy can be visualized by the picture of wavefront in the proper orientation with approximate angle extension (blurring) given by $\Delta\theta\simeq v_{z}(v_{y}/c^{2})$. This relation specifies the limits within which the non relativistic theory can be applied. In fact, it follows that for a very non relativistic electron beam for which $v_{z}^{2}/c^{2}$ is very small, the angle ”blurring” becomes very small too. In this case angle of wavefront tilt $\theta=v_{y}/v_{z}$ is practically sharp $\Delta\theta/\theta\simeq v_{z}^{2}/c^{2}\ll 1$. This is a limiting case of non-relativistic kinematics. The angle ”blurring” is a peculiarity of relativistic beam motion. In the ultrarelativistic limit when $v_{z}\simeq c$, the wavefront tilt has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this tilt could be ascertained.

### v10

@@ 19.6.3 On the Advanced ”Paradox” Related to the Coupling Fields and Particles
Now we consider another situation. Let us see what happens with a kicker, which is installed in the straight section upstream of the bending magnet and is characterized by a kick angle $\theta_{k}=\theta_{0}$. When the kick in the $y$ direction is introduced, there is a redshift of the critical wavelength which arises because, according to corrected addition velocities law, the electron velocity decreases from $v$ to $v-v\theta_{0}^{2}/2$ after the kick. The red shift of the critical frequency $\omega_{c}$ can be expressed by the formula $\Delta\omega_{c}/\omega_{c}=-(3/2)\gamma^{2}\theta_{0}^{2}$. We see a second order correction $\theta_{0}^{2}$ that is, however, multiplied by a large factor $\gamma^{2}$. The result of the covariant approach depends on the absolute value of the kick angle $\theta_{0}$ and the radiation along the velocity direction has a redshift only when the kick angle has a nonzero value. The implicit ”absolute” acceleration means acceleration relative to the fixed stars.
@@
The difference between these two situations, ending with a final uniform translation along the direction of the magnetic field is very interesting. It comes about as the result of the difference between two Lorentz coordinate systems in the lab frame. By trying to accelerate the electron upstream of the bending magnet we have changed Lorentz coordinates for that particular source. We know that in order to keep a Lorentz coordinates system in the lab frame after the kick we need to perform a clock resynchronization. So we should expect the electron velocity to be changed. The difference between the two setups is understandable. When we do not perturb the electron motion (relative to the fixed stars) upstream of the bending magnet, no clock resynchronization takes place, while when we do perturb the motion, clock resynchronization is introduced.
@@
Figure 59: Two sets of initial conditions result in the same uniform motion along the magnetic field direction in the case of absolute time coordinatization in the lab frame. The magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups.
@@
We would now like to describe an apparent paradox. The argument that the difference between these two situations, ending with a final uniform translation along the magnetic field direction, is paradoxical can be summarized in the following way: in the case of absolute time coordinatization in the lab frame, the initial conditions at the bending magnet entrance are apparently identical. In fact, the magnitude of the electron velocity and the orientation of the velocity vector with respect to the magnetic field vector are identical in both setups. We must conclude that when we accelerate the electron in the lab frame upstream of the bending magnet, the information about this acceleration is not included in the noncovariant trajectory. Where is the information about the electron acceleration recorded in the case of absolute time coordinatization? Since an electron is a structureless particle, the situation seems indeed paradoxical.
@@
When the electron passes through a kicker, its fields are perturbed, and now include information about the acceleration. According to the old kinematics, the orientation of the virtual radiation phase front is unvaried. However, Maxwell’s equations do not remain invariant with respect to Galilean transformations and, as discussed throughout this book, the choice of the old kinematics implies using anisotropic field equations. As a result, the phase front remains plane but the direction of propagation is not perpendicular to the phase front. In other words, the radiation beam motion and the radiation phase front normal have different directions. Then, having chosen the absolute time synchronization, electrodynamics predicts that the virtual radiation beam propagates in the kicked direction with the phase front tilt $\theta_{k}$. The information about the electron acceleration is recorded in the perturbation of the self-electromagnetic fields of the electron.
@@ 20.1 Introductory Remarks
In the previous chapter, we attempted to answer the question of why the error in radiation theory should have so long remained undetected. According to the covariant approach, the various relativistic kinematics effects concerning the synchrotron radiation setup, turn up in successive orders of approximation. Instead of small (total) velocity parameter $(v/c)$ in the non-relativistic case, we use a small transverse velocity parameter $(v_{\perp}/c)$. In our previous discussion of bending magnet radiation, we learned that a motion of the single ultrarelativistic electron in a constant magnetic field, according to the theory of relativity, influences the kinematics terms of the second order $(v_{\perp}/c)^{2}$ only. It is demonstrated that due to a combination of the ultrarelativistic (i.e. paraxial) approximation and a very special symmetry of the conventional synchrotron radiation setup, there is a cancellation of the second-order relativistic kinematics effects except the non-zero redshift of the critical frequency, which arises when there are perturbations of the electron motion in the (bending) magnetic field direction. However, synchrotron radiation from bending magnets is emitted within a wide range of frequencies, and the output intensity is not sensitive on the redshift. That means that the spontaneous synchrotron radiation does not show sensitivity to the difference between covariant and non-covariant particle trajectories.
@@
But in the 21st century with the operation XFELs, this situation changes. An XFEL is an example where the first-order kinematics term $(v_{\perp}/c)$ plays an essential role in the description of the XFEL radiation and, in this case, covariant coupling of fields and particles predicts an effect in complete contrast to the conventional treatment. In this chapter, we present a critical reexamination of existing XFEL theory. The main emphasis of this chapter is on coherent undulator radiation from the modulated electron beam. This chapter mainly addressed to readers with limiting knowledge of accelerator and XFEL physics.
@@
Figure 60: A well-known result of conventional (non-covariant) particle tracking. A micro-bunching electron beam passes through a weak dipole magnet (kicker) and undergoes a kick of an angle $\theta_{k}$. The propagation axis of the electron beam is deflected, while the wavefront orientation is preserved.
@@
The usual XFEL theory is based on the use of old Newtonian kinematics for particle dynamics and Einstein’s kinematics for electrodynamics. In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities. For the rectilinear motion of the modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking. However, one of the consequences of the relativity of simultaneity (i.e. mixture of positions and time) is a difference between covariant and non-covariant kinematics of a modulated electron beam in a given magnetic field. The theory of relativity shows that the discussed above difference is related to the acceleration along curved trajectories.
There are several cases where the first-order relativistic effect can occur in XFELs, mainly through the introduction of a trajectory kick  [^68] An angular kick is often an essential part of many XFEL-related diagnostic or experimental procedures. The standard gain length measurement procedure in XFELs employs such kicks. Other applications include ”beam-splitting” schemes where different polarization components are separated by means of an angular kick to the modulated electron beam [73, 74]..
@@
The most elementary of the effects that represent a crucial test of the correct coupling fields and particles is a problem involving the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator.
@@
It would be well to begin with a bird’s view of some of the main results. Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion. A well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before (Fig. 60). In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e. in a wavefront tilt). In existing XFEL theory the wavefront tilt is considered as real. According to conventional treatment, a transverse kick does not change the orientation of a modulation wavefront and hence suppresses the radiation emitted in the direction of the electron motion  [^69] In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis. All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis. However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”). The discrepancy between the directions of the electron motion and wavefront normal after the kick has been discussed in the literature. It was pointed out that coherent radiation is emitted towards the modulation wavefront normal. Thus, according to the conventional coupling of fields and particles (which we claimed was incorrect), the discrepancy between the two directions decreases the radiation efficiency [75]. Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements. The pleasant surprise is that the tolerances predicted are more stringent than they need to be according to the corrected (4D) XFEL theory. This can be considered as one of the reasons for the exceptional progress in XFEL developments over the last decades..
@@
Figure 61: A result of covariant particle tracking. In the ultrarelativistic limit, the orientation of the modulation wavefront, i.e. the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulation electron beam is treated using Lorentz coordinates. The theory of relativity dictates that a modulated electron beam in the ultrarelativistic limit has the same kinematics, in Lorentz coordinates, as a laser beam. According to Maxwell’s equations, the wavefront of a laser beam is always orthogonal to the propagation direction.
@@
The covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment. Namely, in the ultrarelativistic limit, the wavefront of modulation, a plane of simultaneity, is always perpendicular to the electron beam velocity (Fig. 61). As a result, Maxwell’s equations predict a strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction. Experiments show that this prediction is, in fact, true. The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory, see Section 20.6 for more details.
@@
It is worth remarking that the absence of a dynamical explanation for the modulation wavefront readjusting in the Lorentz coordinatization has disturbed some XFEL experts. We only wish to emphasize that a good way to think of the modulation wavefront readjusting is to regard it as a result of the transformation to a new time variable in the framework of the Galilean (”single frame”) electrodynamics.
@@ 20.2 Modulation Wavefront Orientation
Let us suppose that a modulated electron beam moves along the $z$-axis of a Cartesian $(x,y,z)$ system in the lab frame. As an example, suppose that the modulation wavefront is perpendicular to the velocity $v_{z}$. How to measure this orientation? A moving electron bunch changes its position with time. The natural way to do this is to answer the question: when does each electron cross the $y$-axis of the reference system? If we have adopted a method for timing distant events (i.e. a synchronization convention), we can also specify a method for measuring the orientation of the modulation wavefront: if electrons located at the position with maximum density cross the $y$-axis simultaneously at a certain position $z$, then the modulation wavefront is perpendicular to $z$-axis. In other words, the modulation wavefront is defined as a plane of simultaneous events (the events being the arrival of particles located at maximum density): in short, a ”plane of simultaneity”.
@@
Let us formulate the initial conditions in the lab frame in terms of the orientation of the modulation wavefront and beam velocity. Suppose that $v_{z}$ is the velocity of the comoving frame $R(\tau)$ with respect to the lab frame $K(\tau)$ along the common $z$-axis in a positive direction. In the lab frame, we select a special type of coordinate system, a Lorentz coordinate system to be precise. Within a Lorentz frame (i.e. inertial frame with Lorentz coordinates), Einstein’s synchronization of distant clocks and Cartesian space coordinates $(x,y,z)$ are enforced. To have this, we impose that $R$ is connected to $K$ by the Lorentz boost $L(\vec{v}_{z})$, with $\vec{v}_{z}$, which transforms a given four vector event $X$ in space-time into $X_{R}=L(\vec{v_{z}})X$.
@@
We now consider the acceleration of the beam in the lab frame up to velocity $v_{y}$ along the $y$-axis. The question arises of how to assign synchronization in the lab frame after the beam acceleration. Before acceleration, we picked a Lorentz coordinate system. Then, after the acceleration, the beam velocity changes of a small value $v_{y}$ along the $y$-axis. Without changing synchronization in the lab frame after the particle acceleration we have a complicated situation concerning the electrodynamics of moving charges. As a result of such a boost, the transformation of time and spatial coordinates has the form of a Galilean transformation. To keep a Lorentz coordinate system in the lab frame after acceleration, one needs to perform a clock resynchronization by introducing the time shift $t\to t+yv_{y}/c^{2}$. This form of time transformation is justified by the fact that we are dealing with a first-order approximation. Therefore, $v_{y}/c$ is so small that $v_{y}^{2}/c^{2}$ can be neglected and one arrives at the coordinate transformation $y\to y+v_{y}t$, $t\to t+yv_{y}/c^{2}$. The Lorentz transformation just described differs from a Galilean transformation just by the inclusion of the relativity of simultaneity, which is the only relativistic effect that appears in the first order in $v_{y}/c$. The relation $X_{R}=L(\vec{v_{z}})L(\vec{v_{y}})X$ presents a step-by-step change from the lab reference frame $K(\tau+d\tau)$ to $K(\tau)$ and then to the proper reference frame $R$. The shift in the time when electrons located at the position with maximum density cross the $y$-axis of the lab frame $\Delta t=yv_{y}/c^{2}$ has the effect of a rotation of the modulation wavefront on the angle $v_{z}\Delta t/y=v_{z}v_{y}/c^{2}$ in the first order approximation. In ultrarelativistic limits, $v_{z}\simeq c$, and the modulation wavefront rotates exactly as the velocity vector $\vec{v}$.
@@
What does this wavefront readjustment mean in terms of measurements? In absolute time coordinatization the simultaneity of a pair of events has absolute character. The absolute character of the temporal coincidence of two events is a consequence of the absolute time synchronization convention. According to this old kinematics, the modulation wavefront remains unvaried. However, according to the covariant approach, we establish a criterion for the simultaneity of events, which is based on the invariance of the speed of light. It is immediately understood that, as a result of the motion of electrons along the $y$ axis (i.e. along the plane of simultaneity before the boost) with the velocity $v_{y}$, the simultaneity of different events is no longer absolute, i.e. independent of the kick angle $\theta=v_{y}/c$. This reasoning is in analogy with Einstein’s train-embankment thought experiment.
@@
The wavefront orientation has no exact objective meaning because the relativity of simultaneity takes place. The statement that the wavefront orientation has objective meaning within a certain accuracy can be visualized by the picture of the wavefront in the proper orientation with approximate angle extension (blurring) given by $\Delta\theta\simeq v_{z}(v_{y}/c^{2})$. This relation specifies the limits within which the nonrelativistic theory can be applied. It follows that for a very nonrelativistic electron beam for which $v_{z}^{2}/c^{2}$ is very small, the angle ”blurring” becomes very small too. In this case angle of wavefront tilt $\theta=v_{y}/v_{z}$ is practically sharp $\Delta\theta/\theta\simeq v_{z}^{2}/c^{2}\ll 1$. This is a limiting case of non-relativistic kinematics. The angle ”blurring” is a peculiarity of relativistic beam motion. In the ultrarelativistic limit when $v_{z}\simeq c$, the wavefront tilt has no exact objective meaning at all since, due to the finiteness of the speed of light, we cannot specify any experimental method by which this tilt could be ascertained.